If you deposit $20 monthly into an account that pays 12% annually (but compounds the interest monthly), what is the value of the account after 1 year? Answer in dollars and cents
Answer #1

This question requires application of FV of annuity formula, according to which

r = 12%/12 = 1% (monthly), n = 1 * 12 months = 12

FV = 20 * [\frac{(1 + 0.01)^{12} - 1}{0.01}]

FV = 20 * 12.6825

FV = $253.65